# Directionality of the Voynich Script

### v3

@@ Introduction
While the Voynich Manuscript was almost certainly written left-to-right (LTR), the question whether it reads LTR or right-to-left (RTL) has received little quantitative attention. We introduce a statistical method that leverages n-gram perplexity asymmetry to determine directional bias in character sequences. This language-agnostic method computes n-gram perplexities for text read both LTR and RTL. The difference $\Delta=X_{\text{LTR}}-X_{\text{RTL}}$ of cross-entropy $X$ indicates directional optimization:
@@
We selected 3 corpora in our study: the RF1b-e EVA transcription of the Manuscript [4, 5], and two reference data sets for directionality control:
  • LTR control: Melville’s Moby Dick [1]
  • RTL control: The SVLM Hebrew Wikipedia Corpus [2]
All three corpora were preprocessed and tokenized consistently to ensure a fair cross-linguistic comparison.
To normalize sources of directional signal, we implemented two tokenization modes:
  • baseline: tokenization preserves the original grapheme order. Sentences are flattened to a sequence of graphemes (no artificial word markers). We used this mode for Voynich and English tokenization.
  • visual: reverse graphemes inside each word and then reverse the order of words. This models the surface visual order of RTL scripts (what a reader sees on a page rendered RTL). We used this mode for Hebrew tokenization.
@@
Perplexity computation.
@@
Applying this analysis to the Voynich manuscript:
  • $n=2$: $\Delta\approx 0.0653$ (RTL lower perplexity) with 95% CI $(0.0630,0.0677)$
@@
Comparison to English
  • $n=2$: $\Delta\approx-0.0296$ (LTR lower perplexity) with 95% CI (0.0305,0.0288)(-0.0305,-0.0288)
  • $n=3$: $\Delta\approx-0.0132$ (LTR lower) with 95% CI (0.0137,0.0127)(-0.0137,-0.0127)
For English, $\Delta<0$, confirming LTR optimization.
Comparison to Hebrew
  • $n=2$: $\Delta\approx 0.0460$ (RTL lower perplexity) with 95% CI $(0.0454,0.0465)$
@@
For Hebrew, $\Delta$ shifts from positive to negative depending on $n$ due to morphological structure. While mixed results are obtained here, Hebrew’s inconsistency demonstrates that our method is sensitive:
@@
  • Voynich’s consistent RTL pattern stands out more clearly against Hebrew’s complexity
@@
Perplexity-based $\Delta$ captures sequential directional dependencies across the token stream, which are insensitive to local manipulations of word-initial or word-final grapheme distributions.
Words symmetry and boundary diagnostics
The words boundary hypothesis
Universality of the hypothesis
The Voynich script breaks the words boundary hypothesis.
Figure 1: English plateau distribution
Figure 2: Hebrew plateau distribution
Figure 3: Voynich Zipf distribution
@@ Conclusion
Voynich shows consistent positive $\Delta$ (RTL lower perplexity) at $n=2$ and $n=3$ with tight bootstrap CIs. Although we cannot exclude the RTL signal could be a byproduct of a cipher or a scribal artifact rather than a natural language, this indicates local n-gram chains are directionally structured in RTL even while boundary asymmetries are suppressed.
This finding:
  • Supports the hypothesis of RTL reading, independent of visible ink direction.
@@
We also demonstrate that conventional Gini and entropy metrics are unsuitable for measuring directionality in texts with Zipfian character distributions like Voynich. More generally, such unique distribution may explain why conventional linguistic approaches have struggled with the Voynich text.
@@
Future Work
  • For Cipher Theories: Our findings suggest complex cipher models (like homophonic ciphers) over simple substitution ciphers. Future cipher hypothesis should consider not only decode the text but also explain why the encryption process resulted in this specific statistical artifact (Zipfian boundaries) rather than the expected natural language plateau.
  • For Language Theories: These findings suggest that viable language theories may need to account for a natural language whose morphology departs from a plateau boundary distribution. This makes theories involving a constructed language (with artificial, repetitive rules) more plausible.
  • For Analytical Methods: These results suggest that new, customized analytical methods are needed. A promising avenue is to develop models tailored for word boundaries, and to explore the structural behavior of graphemes through n-gram relationships.
Appendix A: Methodological Validation with Kneser-Ney Smoothing
@@
To validate that our directional signals reflect genuine sequential structure rather than statistical artifacts, we implemented a randomization control. For each sentence in our corpora, we randomly permute the token order within sentences while preserving sentence boundaries and special markers. This destroys any genuine directional structure while maintaining vocabulary, sentence lengths, and overall statistical properties.

### v4

@@ Introduction
While the Voynich Manuscript was almost certainly written left-to-right (LTR), the question of whether it should be read LTR or right-to-left (RTL) has received little quantitative attention. We introduce a statistical method that leverages n-gram perplexity asymmetry to determine directional bias in character sequences. This language-agnostic method computes n-gram perplexities for text read both LTR and RTL. The difference $\Delta=X_{\text{LTR}}-X_{\text{RTL}}$ of cross-entropy $X$ indicates directional optimization:
@@
We select five corpora in our study: the RF1b-e EVA transcription of the Manuscript [6, 7], and four reference data sets for directionality control:
  • LTR control (English and French): Melville’s Moby Dick [1], Dumas’ Comte de Monte Cristo [2]
  • RTL control (Hebrew and Arabic): The Big Arabic Corpus [3], The SVLM Hebrew Wikipedia Corpus [4]
All corpora are preprocessed and tokenized consistently to ensure a fair cross-linguistic comparison. To normalize sources of directional signal, we implement two tokenization modes:
  • baseline: tokenization preserves the original grapheme order. Sentences are flattened to a sequence of graphemes (no artificial word markers). We use this mode for Voynich, French and English tokenization.
  • visual: reverse graphemes inside each word and then reverse the order of words. This models the surface visual order of RTL scripts (what a reader sees on a page rendered RTL). We use this mode for Arabic and Hebrew tokenization.
@@
We use two smoothing methods for perplexity computation: Laplace and Kneser-Ney. Both give very similar outcomes.
Laplace smoothing
@@
Kneser-Ney smoothing
Please refer to appendix A for computation and detailed results.
@@
Applying Laplace smoothing perplexity to the Voynich manuscript:
  • $n=2$: $\Delta\approx 0.0653$ (RTL lower perplexity) with 95% CI $(0.0631,0.0677)$
@@
Comparison to LTR languages
For English:
  • $n=2$: $\Delta\approx-0.0296$ (LTR lower perplexity) with 95% CI (0.0306,0.0287)(-0.0306,-0.0287)
  • $n=3$: $\Delta\approx-0.0132$ (LTR lower) with 95% CI (0.0136,0.0127)(-0.0136,-0.0127)
  • $n=4$: $\Delta\approx-0.0033$ (LTR lower) with 95% CI (0.0034,0.0032)(-0.0034,-0.0032)
For French:
  • $n=2$: $\Delta\approx-0.1197$ (LTR lower perplexity) with 95% CI (0.1210,0.1185)(-0.1210,-0.1185)
  • $n=3$: $\Delta\approx-0.0448$ (LTR lower) with 95% CI (0.0454,0.0444)(-0.0454,-0.0444)
  • $n=4$: $\Delta\approx-0.0068$ (LTR lower) with 95% CI (0.0069,0.0067)(-0.0069,-0.0067)
In both languagres, $\Delta<0$, confirming LTR optimization.
Comparison to RTL languages
For Arabic:
  • $n=2$: $\Delta\approx-0.0359$ (LTR lower perplexity) with 95% CI (0.0376,0.0342(-0.0376,-0.0342
  • $n=3$: $\Delta\approx 0.0068$ (RTL lower) with 95% CI $(0.0063,0.0072)$
  • $n=4$: $\Delta\approx 0.0003$ (RTL lower) with 95% CI $(0.0002,0.0003)$
For Hebrew:
  • $n=2$: $\Delta\approx 0.0460$ (RTL lower perplexity) with 95% CI $(0.0454,0.0466)$
@@
  • $n=4$: $\Delta\approx-0.0020$ (LTR lower) with 95% CI (0.0020,0.0020)(-0.0020,-0.0020)
For RTL languages, $\Delta$ shifts from positive to negative depending on $n$ due to morphological structure. While mixed results are obtained here, the shifts demonstrate that our method is sensitive:
@@
  • Voynich’s consistent RTL pattern stands out more clearly against semitic languages’ complexity
@@
Perplexity-based $\Delta$ with Laplace smoothing captures sequential directional dependencies across the token stream, which are insensitive to local manipulations of word-initial or word-final grapheme distributions typically found in natural language directionality studies (see appendix).
@@ Conclusion
Voynich shows consistent positive $\Delta$ (RTL lower perplexity) at $n=2$ to $n=4$ with tight bootstrap CIs. Although we cannot exclude the RTL signal could be a byproduct of a cipher or a scribal artifact rather than a natural language, this indicates local n-gram chains are directionally structured in RTL even while boundary asymmetries are suppressed.
The unique combination of low perplexity and high prediction accuracy at 4-grams level demonstrates that the Voynich manuscript has weak but extremely consistent directional bias.
Our research:
  • Supports the hypothesis of RTL reading if Voynich is a natural language, independent of visible ink direction.
@@
We also demonstrate that conventional Gini and entropy metrics are unsuitable for measuring directionality in texts with Zipfian character distributions like Voynich.
@@
Discussion And Future Work
The exceptional directional predictability of the Voynich manuscript raises profound questions about its nature. It supports several theoretical frameworks:
Complex Encoding Hypothesis
The results are consistent with a sophisticated cipher or encoding scheme that preserves or even amplifies directional structure. Simple substitution ciphers would not maintain the higher-order dependencies observed, suggesting more complex cryptographic methods.
Artificial Language Construction
Viable language theories may need to account for a natural language whose word boundaries departs from a plateau boundary distribution. This makes theories involving a constructed language (with artificial, repetitive rules) more plausible.
The systematically enhanced directional patterns could reflect deliberate linguistic engineering, where directional rules were explicitly designed rather than naturally evolved.
Unique Statistical Properties
The Voynich manuscript may represent a fundamentally different type of text with statistical properties unlike any known natural language, potentially explaining centuries of failed decipherment attempts. Our results suggest that new, customized analytical methods are needed. A promising avenue is to develop models tailored for word boundaries, and to explore the structural behavior of graphemes through n-gram relationships.
Appendix A: Additional Validation Steps
Kneser-Ney Smoothing
@@
To validate that our directional signals reflect genuine sequential structure rather than statistical artifacts, we implement a randomization control. For each sentence in our corpora, we randomly permute the token order within sentences while preserving sentence boundaries and special markers. This destroys any genuine directional structure while maintaining vocabulary, sentence lengths, and overall statistical properties.
@@ Results
Predictive Validation of Directional Analysis
While the perplexity asymmetry method provides evidence for directional optimization, a critical question remains: can the observed patterns generalize to unseen text? To address this, we developed a predictive validation approach that transforms directional analysis from descriptive statistics to testable predictions.
Methodology
We extend the standard perplexity asymmetry analysis with a train-test validation framework:
  1. 1.
    Data Splitting: For each corpus, we randomly partition sentences into 80% training and 20% test sets, maintaining the same split used in the original perplexity analysis.
  2. 2.
    Model Training: Using the training set, we construct separate $n$-gram language models for both reading directions:
    • $M_{\text{LTR}}$: trained on sentences in their original left-to-right order
    • $M_{\text{RTL}}$: trained on sentences with reversed token order
  3. 3.
    Prediction: The cross-entropy difference $\Delta$ provides a corpus-level measure of directional bias. To transform this into a sentence-level predictive task, we use the underlying log-probabilities. For each test sentence $s$, we compute log-probabilities under both models:
    $\ell_{\text{LTR}}(s)$ $=\log P(s|M_{\text{LTR}})$ (1)
    $\ell_{\text{RTL}}(s)$ $=\log P(\text{reverse}(s)|M_{\text{RTL}})$ (2)
    Note that a model with lower cross-entropy (indicating better predictive performance) will necessarily assign a higher log-probability to the data:
    $\hat{d}(s)=\begin{cases}\text{LTR}&\text{if }\ell_{\text{LTR}}(s)>\ell_{\text{RTL}}(s)\\ \text{RTL}&\text{otherwise}\end{cases}$ (3)
  4. 4.
    Evaluation: We compare predictions against ground truth orientations (LTR for English and French, RTL for Hebrew and Arabic) and compute classification accuracy. In the case of EVA, given our results on perplexity, we consider RTL to be the gold orientation.
This approach tests whether directional patterns learned from training data can successfully predict the reading direction of unseen sentences, providing stronger evidence for genuine directional structure than perplexity differences alone.
Results
Table 4 presents the predictive validation results across all tested corpora and $n$-gram orders.
Table 4: Predictive validation results for directional analysis across languages and $n$-gram orders. Accuracy represents the fraction of test sentences where the predicted reading direction matches the ground truth.
Language Test Sentences n=2 $\Delta$ n=2 Accuracy n=3 $\Delta$ n=3 Accuracy n=4 $\Delta$ n=4 Accuracy
English 42,941 0.030-0.030 0.521 0.013-0.013 0.431 0.003-0.003 0.382
French 27,651 0.120-0.120 0.689 0.045-0.045 0.461 0.007-0.007 0.348
Hebrew 77,662 +0.046+0.046 0.610 0.027-0.027 0.519 0.002-0.002 0.610
Arabic 15,694 0.036-0.036 0.429 +0.007+0.007 0.509 +0.000+0.000 0.648
Voynich (EVA) 7,404 +0.065 0.621 +0.017 0.614 +0.006 0.739
Analysis
Voynich Exhibits Exceptional Directional Predictability
The Voynich manuscript achieves the highest predictive accuracy (73.9% at $n=4$), significantly outperforming all natural languages tested. This represents a 47.8% improvement over random guessing and exceeds the performance of known RTL languages Hebrew (61.0%) and Arabic (64.8%).
This pattern is striking when compared to the very low $\Delta$ observed in 4-grams perplexity. It suggests the Voynich text has a systematic, pervasive directional structure that affects sentences in the same subtle way, rather than having some sentences with strong RTL bias and others neutral.
Higher-Order Dependencies Strengthen Voynich Directionality
Unlike natural languages, which typically show degraded performance at higher $n$-gram orders, the Voynich manuscript’s directional predictability increases with context length. The accuracy progression (62.1% → 61.4% → 73.9%) suggests that longer-range dependencies carry stronger directional information than local bigram patterns.
Long-range dependencies are especially valuable since they didn’t surface in our perplexity analysis.
Natural Language Performance Patterns
Natural languages show varied and generally weaker predictive performance:
  • French achieves strong bigram performance (68.9%) but rapidly degrades at higher orders
  • English shows consistently weak directional signals across all $n$-gram lengths
  • Hebrew and Arabic display moderate but inconsistent directional predictability
Methodological Implications
The predictive validation provides several advantages over traditional perplexity asymmetry analysis:
  • Generalization Testing: Demonstrates that directional patterns extend beyond training data
  • Effect Size Quantification: Accuracy metrics are more interpretable than cross-entropy differences, and show long-range dependencies
  • Artifact Detection: Random text or methodological artifacts would yield  50% accuracy
Appendix B: Words symmetry and boundary diagnostics
The words boundary hypothesis
Universality of the hypothesis
The Voynich script breaks the words boundary hypothesis.
Figure 1: English plateau distribution
Figure 2: Hebrew plateau distribution
Figure 3: Voynich Zipf distribution